I am SO STUCK on this problem... PLEASE HELP ASAP!!!
Suppose Kevin and Jill both deposit $4000 into their personal accounts. If Kevin’s account earns 5% simple interest annually and Jill’s earns 5% interest compounded annually, how much will each account balance show at the end of 5 years? Calculate the difference between each account.

Answer by mananth(16946)

KEVINS ACCOUNT
p= 4000
R= 0.05
T= 5
iNTEREST = 4000*0.05*5=1000
Kevin's account has $ 5000
-----
JILLS account
P =4000
r=0.05
n=1
t=5
Principal P = 4000
Amount= A
years=n 5.00
compounded 1 times a year t
Rate = 5.00 0.05
Amount = P*((n+r)/n)^n*t

Amount =4000*(1 +0.05/t)^5*1
Amount =4000*(1+0.05)^5
4000 *( 1.05 )^ 5
Amount = 5105.13
you can work out the difference